OpenStudy (anonymous):

find interval of convergence for x^((n^2+2)/(n+3)), n=1->∞

OpenStudy (mr.math):

Is it \(\large x^{\frac{n^2+2}{n+3}}\)?

OpenStudy (mr.math):

If so, then you can use the ratio test. \[\large |{a_{n+1} \over a_n}|= \large |{x^{n^2+2n+3 \over n+4} \over x^{n^2+3 \over n+3}}|\] \[{n^2+2n+3 \over n+4}-{n^2+2 \over n+3}=({11 \over n+4}-{11 \over n+3}+1)\to 1 \text{ as } n \to \infty\] Thus \(\large \lim_{n\to \infty} |{a_{n+1} \over a_n}|=x.\)

OpenStudy (mr.math):

That means that the series converges for \(-1<x<1\). You need now to test the end points -1 and 1.

OpenStudy (anonymous):

i got the same result, but in my book says: absolute convergent for 0<x<1, divergent for x>=1

OpenStudy (asnaseer):

I would suggest that your book is wrong - @Mr.Math is definitely correct here. @Mr.Math made just one "typo" error where he has \(x^{\frac{n^2+3}{n+3}}\) in the denominator of the ratio test instead of \(x^{\frac{n^2+2}{n+3}}\) - i.e. should be \(n^2+2\) not \(n^2+3\), but that error was corrected in the next line.